OpenStudy (idealist10):

A stone weighing 1/2 lb is thrown upward from an initial height of 5 ft with an initial speed of 32 ft/s. Air resistance is proportional to speed, with k=1/128 lb-s/ft. Find the maximum height attained by the stone.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

 As before, we have the same setup: \[\begin{align*} mg+kv&=m\frac{dv}{dt}\\ \left(\frac{1}{2}\right)\left(-32\right)-\left(\frac{1}{128}\right)v&=\frac{1}{2}\frac{dv}{dt}\\ \frac{dv}{dt}+\frac{1}{64}v&=-32 \end{align*}\] and you're given the initial velocity, so \(v(0)=32\). Additionally, you're given the initial height, so you can integrate the velocity function and use the initial condition \(h(0)=5\), where \(v(t)=h'(t)\). Alternatively, you can determine when \(v(t)=0\).

OpenStudy (idealist10):

So I got\[v=-2048+2080e ^{-\frac{ t }{ 64 }}\]

OpenStudy (idealist10):

Since v(t)=h'(t), I integrated v(t) and got \[h(t)=-2048t-133120e ^{-\frac{ t }{ 64 }}+133125\]

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

mg+kv=m(dv/dt) since W=mg=-1/2 because it's downwards, -1/2-(1/128)v=(1/64)(dv/dt) since you said g=32, I got mass=1/64 dv/dt+(1/2)v=-32 integrating factor e^(t/2) e^(t/2)v'+e^(t/2)/2*v=-32e^(t/2) e^(t/2)v=-64e^(t/2)+C v=-64+Ce^(-t/2) v(0)=32 C=98 v=-64+98e^(-t/2) v(t)=h'(t) h'(t)=-64+98e^(-t/2) integrate h'(t) h(t)=-64t-196e^(-t/2)+C h(0)=5 C=201 h(t)=-64t-196e^(-t/2)+201 v(t)=0 when t=0.852169 h(0.852169)=18.4612 But the answer in the book is 17.10 ft. This problem is driving me crazy. What's wrong?

OpenStudy (anonymous):

Oh hold up, looks like \(C=96\), not \(98\).

OpenStudy (anonymous):

\[32=-64+Ce^{-0/2}~~\implies~~96=C\]

OpenStudy (idealist10):

HAHAHAHAHA, I know, right? I'm so sorry about that. Thank you so much for pointing out the error. You're so good.

OpenStudy (idealist10):

YES! I finally got the correct answer! Thank you so much for pointing out the error.

OpenStudy (anonymous):

You're welcome!